OpenStudy (migitmack):

An online movie rental store rented out 6,816 movies one month. This was 40% less than the number of movies it rented out the previous month. The following equation represents this situation, where x represents the number of movies the store rented out the previous month. x − 0.4x = 6,816 How many movies did the store rent out in both months combined?

OpenStudy (anonymous):

40/100 * 6816= 2726.4 last month = 6816+2726=9542.4 this month = 6816 combined = 16 358 sorry if i'm wrong

OpenStudy (migitmack):

ur wrong

OpenStudy (anonymous):

:( okay ..